OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x2-6x+8=y (the x is squared) A. (–3,– 6) B. (3,–1) C. (–3,1) D. (–1, 3)

OpenStudy (chaise):

(x-h)^2+k has a vertex of (h,k) x^2-6x+8=y x2-6x =y-8 x^2-6x+9=y-8+9 (x-3)^2=y+1 (x-3)^2-1=y Vertex is (3,-1)